# FSTA-SNN:Frequency-based Spatial-Temporal Attention Module for Spiking Neural Networks

### v1

@@ Abstract
Spiking Neural Networks (SNNs) are emerging as a promising alternative to Artificial Neural Networks (ANNs) due to their inherent energy efficiency. Owing to the inherent sparsity in spike generation within SNNs, the in-depth analysis and optimization of intermediate output spikes are often neglected. This oversight significantly restricts the inherent energy efficiency of SNNs and diminishes their advantages in spatiotemporal feature extraction, resulting in a lack of accuracy and unnecessary energy expenditure. In this work, we analyze the inherent spiking characteristics of SNNs from both temporal and spatial perspectives. In terms of spatial analysis, we find that shallow layers tend to focus on learning vertical variations, while deeper layers gradually learn horizontal variations of features. Regarding temporal analysis, we observe that there is not a significant difference in feature learning across different time steps. This suggests that increasing the time steps has limited effect on feature learning. Based on the insights derived from these analyses, we propose a Frequency-based Spatial-Temporal Attention (FSTA) module to enhance feature learning in SNNs. This module aims to improve the feature learning capabilities by suppressing redundant spike features. The experimental results indicate that the introduction of the FSTA module significantly reduces the spike firing rate of SNNs, demonstrating superior performance compared to state-of-the-art baselines across multiple datasets. Our source code is available in https://github.com/yukairong/FSTA-SNN.
@@ Introduction
Figure 1: Comparison of the output spike frequency spectrum distribution of SNNs across different model structures, datasets, layer depths, and temporal perspectives. (a) The spectral distribution of spike outputs at different layers of the ResNet architecture at the same time step in static dataset. (b) The spectral distribution of spike outputs at different layers of the ResNet architecture at the same time step in dynamic dataset. (c) The spectral distribution of spike outputs at different layers of the VGG architecture at the same time step in static dataset. (d) The temporal variation of the frequency bands where spectral energy is most concentrated across different layers.
@@
  • We conduct an in-depth study of SNN learning preferences, presenting a novel frequency-based analysis of spikes for the first time. This perspective is crucial for improving sparsification and energy efficiency in neuromorphic computing, while also providing a theoretical basis for enhancing SNN performance.
  • We propose a Frequency-based Spatial-Temporal Attention (FSTA) module, a plug-and-play component that introduces a minimal number of additional parameters. This module effectively reduces spike firing rate while boosting performance, leading to lower energy consumption and improved computational efficiency.
@@
We also analyze the spike spectral distribution of the model using DFT with the proposed method. As shown in Fig. 5, the frequency distribution remains largely consistent with vanilla SNN after incorporating the FSTA module; However, the magnitudes have significantly decreased. This reduction indicates that the proposed method effectively suppresses unnecessary spikes, consistent with the reduced spike firing rate reported in Fig. 4. Fig. 5(b) further illustrates that the introduction of the module narrows the region of shallow spectral energy concentration, reducing redundant spike generation. In contrast, deeper layers expand the frequency distribution, revealing new spike features that the vanilla SNN could not capture. The spectral diagrams correspond well with the heat map shown in Fig. 5(c).

### v2

@@ Abstract
Spiking Neural Networks (SNNs) are emerging as a promising alternative to Artificial Neural Networks (ANNs) due to their inherent energy efficiency. Owing to the inherent sparsity in spike generation within SNNs, the in-depth analysis and optimization of intermediate output spikes are often neglected. This oversight significantly restricts the inherent energy efficiency of SNNs and diminishes their advantages in spatiotemporal feature extraction, resulting in a lack of accuracy and unnecessary energy expenditure. In this work, we analyze the inherent spiking characteristics of SNNs from both temporal and spatial perspectives. In terms of spatial analysis, we find that shallow layers tend to focus on learning vertical variations, while deeper layers gradually learn horizontal variations of features. Regarding temporal analysis, we observe that there is not a significant difference in feature learning across different time steps. This suggests that increasing the time steps has limited effect on feature learning. Based on the insights derived from these analyses, we propose a Frequency-based Spatial-Temporal Attention (FSTA) module to enhance feature learning in SNNs. This module aims to improve the feature learning capabilities by suppressing redundant spike features. The experimental results indicate that the introduction of the FSTA module significantly reduces the spike firing rate of SNNs, demonstrating superior performance compared to state-of-the-art baselines across multiple datasets.
Codehttps://github.com/yukairong/FSTA-SNN
@@ Introduction
Figure 1: Comparison of the output spike frequency spectrum distribution of SNNs across different model structures, datasets, layer depths, and temporal perspectives. (a) The spectral distribution of spike outputs at different layers of the ResNet architecture at the same time step in static dataset. (b) The spectral distribution of spike outputs at different layers of the ResNet architecture at the same time step in dynamic dataset. (c) The spectral distribution of spike outputs at different layers of the VGG architecture at the same time step in static dataset. (d) The temporal variation of the frequency bands where spectral energy is most concentrated across different layers. Here, Amp, Freq, and T represent amplitude, frequency range and timesteps.
@@
  • We present a comprehensive study of SNN learning preferences, introducing a novel frequency-based spike analysis. This framework is essential for optimizing sparsification and energy efficiency and provides a theoretical foundation for enhancing SNN performance.
  • We propose a Frequency-Based Spatial-Temporal Attention (FSTA) module, a plug-and-play component that introduces a minimal number of additional parameters. This module effectively reduces spike firing rate while boosting performance.
@@
We also analyze the spike spectral distribution of the model using DFT with the proposed method. As shown in Fig. 5, the frequency distribution remains largely consistent with vanilla SNN after incorporating the FSTA module. However, the magnitudes have significantly decreased. This reduction indicates that the proposed method effectively suppresses unnecessary spikes, consistent with the reduced spike firing rate reported in Fig. 4. Fig. 5(b) further illustrates that the introduction of the module narrows the region of shallow spectral energy concentration, reducing redundant spike generation. In contrast, deeper layers expand the frequency distribution, revealing new spike features that the vanilla SNN could not capture. The spectral diagrams correspond well with the heat map shown in Fig. 5(c).